Rachel’s dad was 30 when she was born now he is 3x older than Rachel. How old is Rachel’s dad?
Found 2 solutions by stanbon, greenestamps:
Answer by stanbon(75887)

Rachel’s dad was 30 when she was born now he is 3x older than Rachel. How old is Rachel’s dad?
------
Equations:
D = R + 30
D = 3R
-----
R + 30 = 3R
2R = 30
R = 15 yrs (Rachel's age now)
D = R + 30 = 45 yrs (Dad's age now)
----------
Cheers,
Stan H.
----------

Answer by greenestamps(13203)



The problem says her dad is "3 times older than" Rachel, meaning the difference in their ages is 3 times Rachel's age. (It does NOT mean that her dad is 3 times AS OLD AS her.)

So the difference in their ages is 3 times Rachel's age; but we know the difference in their ages is 30 years. So Rachel's age is 10; that makes her dad 40.

------------------------------------------------------
Note: The answer by tutor stanbon -- that Rachel's dad is 45 years old -- is PROBABLY the answer the author of the problem wanted.

However, the right answer, using the grammatically correct interpretation of the given information, is that her dad is 40.

The problem is with the phrase saying that her dad is "3x older" than Rachel.

While everyday sloppy usage treats that the same as "3x as old", they mean different things.

For "3x as old", you obviously just multiply by 3.

But the grammatically correct interpretation of "3x older" is that her Dad is her age, plus 3 more times her age -- i.e., 4 times AS OLD.